OpenStudy (anonymous):

Who can help me find area between curves? 7x+y^2=8 x=y i dont know how to approach this one.

OpenStudy (anonymous):

x=1/64 y=1/8 idk if that wat u was lookn for

OpenStudy (anonymous):

i need to find shared area between these 2 curves

OpenStudy (anonymous):

Solve both curves for x, the reason for solving for x is because you don't have to deal with the +/- ambiguity that would come from solving for y. So you have: x=(1/7)(8-y^2) Now, plug in a y value to see which is on the right. In this case, x=y is on the LEFT so you have: (1/7)(8-y^2)=y y^2+7y-8=0 (y+8)(y-1)=0 So your limits of integration are -8 to 1. Your integral then looks like this: \[\int\limits_{-8}^{1} \frac{1}{7}(8-y^2)-ydy=\frac{1}{7}\int\limits_{-8}^{1}8-y^2-7ydy=\frac{1}{7}[8y-\frac{y^3}{3}-\frac{7y^2}{2}]_{-8}^{1}\]

OpenStudy (anonymous):

Which gives you 243/14.

OpenStudy (anonymous):

wait a sec, i'll check it ...

OpenStudy (anonymous):

I just ran it through wolfram.

OpenStudy (anonymous):

thanks, i was confused on the part, where you need to find limits of integration..

OpenStudy (anonymous):

Okay, all I did was set the equations(after solving them for x) equal. So you have x=y; x=(1/7)(8-y^2) Set them equal: y=(1/7)(8-y^2) Multiply by 7 7y=8-y^2 Add the y^2, subtract the 8 y^2+7y-8=0 Factor: (y+8)(y-1)=0 =>y=-8,1

OpenStudy (anonymous):

Is that all you need?

OpenStudy (anonymous):

yea, thanks!

OpenStudy (anonymous):

No problem :P